# Galactic Archaeology with Gaia

### v1

@@ 4.5 The lowest metallicity stars
Stars with the lowest metallicities are likely the oldest formed in a given environment, although the age-metallicity relationships can differ substantially depending on the mass of the host galaxy. Gaia is heavily relied upon to (1) make sense of the previously discovered metal-poor stars, and (2) to identify new low-metallicity stars missed by previous efforts. Before Gaia, no single spectroscopic survey managed to examine the low-metallicity regime of the Galactic halo comprehensively. As a result, the currently available sample of stars with [Fe/H]$<-2$ have strong and heterogeneous selection effects from a patchwork of different studies (e.g. Frebel and Norris, 2015). As the Gaia data started to come out, efforts to consolidate the census of the metal-poor end of the stellar MDF accelerated, as exemplified by e.g. the study of Li et al. (2018). With the DR3, Gaia has stepped up the game by producing the largest homogeneous catalogs of low-metallicity candidates selected using its high-resolution spectroscopy (see e.g. Matsuno et al., 2022; Viswanathan et al., 2023), and low-resolution spectro-photometry (XP spectra, e.g. Andrae et al., 2023; Yao et al., 2024).
@@
The Gaia XP revolution is only beginning, but the first attempts to extract stellar atmosphere parameters and individual abundances from admittedly extremely low-resolution but homogeneous and enormous XP spectra are already bearing fruit. For example, some success has been reported in extracting $\alpha$-abundances from XP (Chandra et al., 2023b). Perhaps, this analysis can even be extended to the low-metallicity regime, where the exact behaviour of the in-situ $\alpha$-abundance remains a puzzle. According to Conroy et al. (2022), stars with [Fe/H]$<-2$ may exhibit noticeably higher $\alpha$ values than previously thought, moving above the established $\alpha$ plateau. While the sensitivity of the XP spectra to changes in $\alpha$-abundance comes as a surprise, spectral variations due to changes in carbon abundance are less subtle and thus can be exploited (see e.g. Witten et al., 2022). To this end, the first all-sky, homogeneous sample of carbon-enhanced metal-poor (CEMP) stars is built by Lucey et al. (2023) using Gaia DR3 XP spectra. CEMPs come in several flavours: those that show signs of $r$ and/or $s$-process enhancement and those that do not (Aoki et al., 2007). The $r/s$-enhanced CEMPs are most common and likely linked to pollution by companions in binary systems (Lucatello et al., 2005; Starkenburg et al., 2014). Instead, CEMP-no stars appear to be mostly single (but see Hansen et al., 2016; Arentsen et al., 2019) and thus their carbon-enhancement is likely primordial, and linked to the very first bouts of star formation in the Universe (Frebel and Norris, 2015). Using Gaia’s astrometry to shed light on the orbital properties of confirmed CEMP stars reveals that many of them are typical halo denizens and can be associated with known halo sub-structures (such as GS/E, see Zepeda et al., 2023).
@@
Plotting global orbital properties of the low-metallicity stars discovered recently reveals a small but noticeable angular momentum asymmetry, with more stars found on prograde orbits at all metallicities, going down as low as [Fe/H]$\approx-6$ (Sestito et al., 2019; Di Matteo et al., 2020). Note, however, that focusing on metal-poor stars with $r$-process enhancement, Roederer et al. (2018) do not see such an angular momentum asymmetry. Many of the low-metallicity stars on prograde orbits also spend most of their time close to the Galactic disc, at low $|z|$. These unexpected trends have inspired several hypotheses, including the suggestion that in the Milky Way, the stellar disc may have formed exceedingly early and at extremely low metallicities (e.g. Di Matteo et al., 2020). Testing this idea against numerical simulations however reveals that in the Milky Way-like progenitors, dominant stellar discs form sufficiently late, at noticeably higher metallicities (e.g. Belokurov and Kravtsov, 2022; Semenov et al., 2023; Chandra et al., 2023a). Sestito et al. (2021) address the origin of the low-metallicity stars on prograde, low-$|z|$ orbits, and conclude that most of them must be accreted. It remains unclear what induced the net spin in the low-metallicity halo, is it inherent or acquired? Dillamore et al. (2023) demonstrate that the Galactic bar can spin up intrinsically non-rotating stellar haloes to velocities similar to those observed. In their models, a halo spun-up by the bar shows a rotation curve declining with Galactocentric radius — this prediction can be tested with upcoming data.
@@ 5.1 Sagittarius and the Milky Way disc
The Sgr dwarf galaxy (Ibata et al., 1994) and its associated stellar stream (e.g. Newberg et al., 2002; Majewski et al., 2003; Belokurov et al., 2006) are an archetype of dwarf galaxy disruption. The stellar stream emanating from a known dwarf galaxy is a clear example of accretion in action, and studies of the stream have been used to measure the Galactic potential as well as the orbit and mass of the Sgr dwarf progenitor (e.g. Law and Majewski, 2010; Dierickx and Loeb, 2017; Gibbons et al., 2017; Fardal et al., 2019; Vasiliev et al., 2021). The mere fact that the progenitor is still intact points to a fairly recent accretion event (since redshift $z\sim 1$), and this has been confirmed by most modeling efforts.
The omniscience of the Sgr stream has been obvious for several years, most notably displayed in the famous ‘field-of-streams’ image from SDSS (Belokurov et al., 2006). However, the Gaia mission has allowed a more detailed examination of this well-known stream (Antoja et al., 2020; Ibata et al., 2020; Ramos et al., 2020). In addition to properties of the stream itself, the Gaia era has provided evidence that Sgr has actually perturbed the Galactic disc during its orbit around the galaxy. Now, this is not a new idea — several past works have shown that a Sgr-like dwarf could induce perturbations to the disc (e.g. Dehnen, 1998; Ibata and Razoumov, 1998; Quillen et al., 2009; Laporte et al., 2018), and there has been evidence of disc perturbations from pre-Gaia data (e.g. Gómez et al., 2012; Widrow et al., 2012; Carlin et al., 2013; Schönrich and Dehnen, 2018). However, Gaia provided unambiguous evidence that the galaxy is out of equilibrium, and is currently undergoing a phase-mixing process from an out-of-equilibrium state. The evidence for this phase-mixing comes from the prominent spiral feature in position and velocity discovered by Antoja et al. (2018) (see Fig. 21). This so-called Gaia phase-space ‘Snail’ potentially arises from phase-mixing of past gravitational disturbances of the Milky Way disc.
@@ 5.2 A massive LMC
The most massive Milky Way satellites, the Large and Small Magellanic Clouds (LMC, SMC), have been known about for centuries. Visible to the naked eye in the Southern Hemisphere, their brightness is unmistakable. However, it is only in recent years that we have accepted the sheer massiveness of the Clouds, particularly the LMC. A simple abundance-matching argument already places the LMC at a significant mass: with a stellar mass of $2.7\times 10^{9}M_{\odot}$ (van der Marel et al., 2002), this roughly corresponds to a halo mass of $10^{11}M_{\odot}$ (Moster et al., 2013). A mass of $\sim 10^{11}M_{\odot}$ for the LMC also agrees with modeling efforts which aim to reproduce the significant 3D velocities of the Clouds and allow for LMC-SMC interactions (e.g. Besla et al., 2010; Kallivayalil et al., 2013). Most tidal models now advocate for a massive satellite (carrying its smaller sibling) that has rapidly joined the Milky Way in the past couple of Gyr (although Vasiliev 2024 show that the observations are also consistent with the LMC being on its second passage around the Milky Way). More recent works using Magellanic ‘satellite-of-satellite’ counts, or modeling of orbital streams in the presence of the LMC now agree that the LMC must have a mass $\gtrsim 1\times 10^{11}M_{\odot}$ (e.g. Erkal et al., 2019; Erkal and Belokurov, 2020)
At the pericenter of the LMC ($\sim 50$ kpc), a mass of $\sim 10^{11}M_{\odot}$ is a significant fraction of the Milky Way halo mass out to that radius (roughly $\sim 1/3$). Gómez et al. (2015) show that a massive LMC at this proximity can significantly displace the center-of-mass (COM) of the Milky Way. Thus, this substantial perturber in our Galaxy cannot be ignored, and the scale of the damage inflicted by the LMC is now starting to be revealed by the Gaia mission.
@@ 5.2.1 LMC wake
Figure 22: Left: A model of the LMC-induced dynamical friction (transient) wake and collective response in the Milky Way halo shown in the Galactocentric YZ plane (Figure 2 in Garavito-Camargo et al., 2021). The colorbar shows the density contrast (light = overdense, dark = underdense). The transient wake follows the past orbit of the LMC (indicated by the red line). The collective response appears predominantly to the North of the Galactic disc (central blue ellipse). Right: Distribution of K giant stars in Milky Way halo at 60 kpc $<R_{\rm gal}<100$ kpc in Galactic coordinates (Figure 1 in Conroy et al., 2021). The overdensity in the North and South-West are attributed to the collective response and local wake, respectively.
@@ 5.2.2 The Milky Way-LMC system
The appreciation of the LMC’s impact on the Milky Way has spurned a new era of dynamical modeling: we now need to consider the ‘Milky Way-LMC system’ (or even the ‘Milky Way-LMC-Sgr system’) rather than just a Milky Way host halo. The fundamental assumptions that govern many modeling techniques, such as equilibrium, static potentials, spherical or axisymmetric symmetry, etc, are no longer valid. What are the implications of this for the many decades of work aiming to measure the total mass of the Galaxy using halo tracer populations (e.g. Wilkinson and Evans, 1999; Xue et al., 2008; Deason et al., 2012; Callingham et al., 2019; Eadie and Jurić, 2019)? A cautionary tale was spun by Erkal et al. (2021) who showed that mass estimates based on equilibrium modeling of tracers that ignore the presence of the LMC are typically biased high, and could even lead to an overestimate of up to 50 percent! On a more positive note, Deason et al. (2021) showed that by allowing for a small perturbation to the velocities of the halo stars a mass profile can be uncovered in a Milky Way-LMC potential. However, this depends on the radial range covered by the data and the region of the sky probed, as these authors focused on a region covered by SDSS where the predicted velocity gradients in the halo are small. Deason et al. (2021) also argued that systematic effects due to substructures, such as shells and clouds, can cause an even bigger effect than the LMC on mass-modeling efforts. Thus, one could argue that despite increasing sample sizes of multi-dimensional data, our aim to measure the Milky Way potential accurately, at least using classical methods, is perhaps waning.
@@
It is worth remarking in the post-Gaia era that we likely need to reassess what the ‘total mass of the Milky Way’ even means. Particularly in light of the messiness caused by the LMC and, to a lesser extent, Sgr. Detailed modeling of the Milky Way-LMC-Sgr system (Vasiliev et al., 2021) has indicated that considering these major components together can prove vital. Indeed, recent stream-modeling techniques have shown that the LMC is required in order to produce stream tracks that are consistent with the data (e.g. Erkal et al., 2019; Vasiliev et al., 2021). However, while tailored models are valuable there are still systematic effects revealed from cosmological simulations of stellar haloes that need to be taken into account in modeling efforts. How do we reconcile these two approaches? Perhaps the most promising avenue is the use of ‘Basis-Function-Expansion’ (BFE) methods (Lowing et al., 2011; Sanders et al., 2020; Garavito-Camargo et al., 2021; Petersen et al., 2022). Here, a time-evolving potential from an $N$-body simulation is approximated using a small number of basis functions. Thus, at a much lower cost than the original simulation, subhalo orbits can be explored in this potential without the need to continually re-simulate the host. In principle, these methods could be used to include the cosmological context of subhalo accretion, whilst retaining the ‘realistic’ potential of the Milky Way-LMC system. For example, Bayesian inference could be applied to a large suite of such models in order to infer global properties of the Galactic potential from the phase-space distribution of halo stars. The BFE models could even provide the framework to combine different dynamical modeling techniques, either from multiple streams and/or halo stars. With the exquisite data at hand, which will only increase in number and quality in the coming years, we clearly need to develop more sophisticated modeling approaches in order to keep pace with the data.
@@ 5.3 Census of accreted substructures
Figure 23: Dynamical substructure identified by Dodd et al. (2023) in the local stellar halo using the Gaia DR3 data. Top: Projections of the energy $E$ and angular momentum $L$ space. Bottom: Projections of the velocity space in cylindrical polars. [From Dodd et al. (2023)].
@@
Figure 24: Streams discovered using STREAMFINDER algorithm, colour-coded according to their metallicity. [From Ibata et al. (2023)].
@@
The identification by Helmi et al. (1999) of a handful of stars belonging to a stellar stream passing through the Solar neighborhood is a good example of the type of archaeological experiment that had appeared trailblazing and challenging before Gaia, but has been made effortless and almost mundane thanks to Gaia today. Since the pioneering study of Helmi et al. (1999) the methods to detect substructure have been sharpened (see e.g. Helmi and de Zeeuw, 2000; Harding et al., 2001; Knebe et al., 2005; Meza et al., 2005; Brown et al., 2005; Bullock and Johnston, 2005; Arifyanto and Fuchs, 2006; McMillan and Binney, 2008; Johnston et al., 2008; Gómez et al., 2010; Bovy et al., 2011; Sanderson et al., 2015; Lisanti et al., 2015; Antoja et al., 2015) whilst the available datasets have grown in size and richness (e.g. Yanny et al., 2000; Vivas et al., 2001; Newberg et al., 2002; Majewski et al., 2003; Clewley and Kinman, 2006; Kepley et al., 2007; Klement et al., 2008; Morrison et al., 2009; Klement et al., 2009; Schlaufman et al., 2009; Smith et al., 2009; Xue et al., 2011; Re Fiorentin et al., 2015; Janesh et al., 2016) but ultimately, the community had to wait for the arrival of the Gaia data to reveal the true wealth of accreted debris. This is particularly true for merger remnants that do not appear spatially coherent, although the detection of long, narrow streams has also evolved to a new level of sophistication thanks to the Gaia data (see e.g. Malhan and Ibata, 2018; Malhan et al., 2018; Ibata et al., 2019, 2023).
@@
Figure 25: Left: Current (black) and initial (red) Globular Cluster masses as a function of their orbital size. Assuming that the GC initial mass function is not a strong function of Galacto-centric radius, a large number of low-mass GC with apocentres inside $\sim$10 kpc appears to have been destroyed by the Galaxy. Middle and Right columns: Current (blue) and initial (red) mass functions of the Milky Way GCs in three different bins of Galacto-centric radius. Two different parametric forms for the initial mass functions are shown: power-law (middle) and log-normal (right). The integral of the difference between the current and the initial mass functions gives an estimate of the total number of GCs destroyed. [Reproduced from Baumgardt et al. (2019)].
@@
Out of the box, the view on the halo substructure with Gaia DR1 appears underwhelming due to the minuscule number of halo stars in the shallow TGAS catalogue (Helmi et al., 2017; Myeong et al., 2017). Using the deeper SDSS-Gaia instead (see also Section 2.1), shows the promise of the modern space astrometry: the halo in the Solar neighborhood is riddled with pieces of accretion events (see Myeong et al., 2018a). This study uncovers $\lesssim 10$ new clumps in velocity space, of which at least one, namely S2, had been known before (Helmi et al., 1999). However, the integrals of motion space are better suited for substructure detection because phase-mixed debris can be smeared into arcs in velocity space; worse still, multiple wraps of the same stream might appear as distinct clumps. The immediate and clear improvement is demonstrated by Myeong et al. (2018b) who use SDSS-Gaia DR1 and translate into the action space, detecting a factor of $\sim 3$ more distinct sub-structures compared to their previous analysis of the same data (see also Wu et al., 2022, for a recent discussion of the pragmatic advantages of using action space). Once the Gaia DR2 was delivered, the hunt for the nearby halo merger remnants begun in earnest (Koppelman et al., 2018; Roederer et al., 2018; Koppelman et al., 2019; O’Hare et al., 2020; Borsato et al., 2020; Necib et al., 2020; Limberg et al., 2021; Lövdal et al., 2022; Tenachi et al., 2022; Ruiz-Lara et al., 2022; Malhan, 2022; Viswanathan et al., 2023; Dodd et al., 2023; Ye et al., 2024).

### v2

@@ 4.5 The lowest metallicity stars
Stars with the lowest metallicities are likely the oldest formed in a given environment, although the age-metallicity relationships can differ substantially depending on the mass of the host galaxy. Gaia is heavily relied upon to (1) make sense of the previously discovered metal-poor stars, and (2) to identify new low-metallicity stars missed by previous efforts. Before Gaia, no single spectroscopic survey managed to examine the low-metallicity regime of the Galactic halo comprehensively. As a result, the currently available sample of stars with [Fe/H]$<-2$ have strong and heterogeneous selection effects from a patchwork of different studies (e.g. Frebel and Norris, 2015). As the Gaia data started to come out, efforts to consolidate the census of the metal-poor end of the stellar MDF accelerated, as exemplified by e.g. the study of Li et al. (2018). With the DR3, Gaia has stepped up the game by producing the largest homogeneous catalogues of low-metallicity candidates selected using its high-resolution spectroscopy (see e.g. Matsuno2022; Viswanathan et al., 2024), and low-resolution spectro-photometry (XP spectra, e.g. Andrae et al., 2023; Yao et al., 2024).
@@
The Gaia XP revolution is only beginning, but the first attempts to extract stellar atmosphere parameters and individual abundances from admittedly extremely low-resolution but homogeneous and enormous XP spectra are already bearing fruit. For example, some success has been reported in extracting $\alpha$-abundances from XP (Li et al., 2024). Perhaps, this analysis can even be extended to the low-metallicity regime, where the exact behaviour of the in-situ $\alpha$-abundance remains a puzzle. According to Conroy et al. (2022), stars with [Fe/H]$<-2$ may exhibit noticeably higher $\alpha$ values than previously thought, moving above the established $\alpha$ plateau. While the sensitivity of the XP spectra to changes in $\alpha$-abundance comes as a surprise, spectral variations due to changes in carbon abundance are less subtle and thus can be exploited (see e.g. Witten et al., 2022). To this end, the first all-sky, homogeneous sample of carbon-enhanced metal-poor (CEMP) stars is built by Lucey et al. (2023) using Gaia DR3 XP spectra. CEMPs come in several flavours: those that show signs of $r$ and/or $s$-process enhancement and those that do not (Aoki et al., 2007). The $r/s$-enhanced CEMPs are most common and likely linked to pollution by companions in binary systems (Lucatello et al., 2005; Starkenburg et al., 2014). Instead, CEMP-no stars appear to be mostly single (but see Hansen et al., 2016; Arentsen et al., 2019) and thus their carbon-enhancement is likely primordial, and linked to the very first bouts of star formation in the Universe (Frebel and Norris, 2015). Using Gaia’s astrometry to shed light on the orbital properties of confirmed CEMP stars reveals that many of them are typical halo denizens and can be associated with known halo sub-structures (such as GS/E, see Zepeda et al., 2023).
@@
Plotting global orbital properties of the low-metallicity stars discovered recently reveals a small but noticeable angular momentum asymmetry, with more stars found on prograde orbits at all metallicities, going down as low as [Fe/H]$\approx-6$ (Sestito et al., 2019; Di Matteo et al., 2020). Note, however, that focusing on metal-poor stars with $r$-process enhancement, Roederer et al. (2018) do not see such an angular momentum asymmetry. Many of the low-metallicity stars on prograde orbits also spend most of their time close to the Galactic disc, at low $|z|$. These unexpected trends have inspired several hypotheses, including the suggestion that in the Milky Way, the stellar disc may have formed exceedingly early and at extremely low metallicities (e.g. Di Matteo et al., 2020). Testing this idea against numerical simulations however reveals that in the Milky Way-like progenitors, dominant stellar discs form sufficiently late, at noticeably higher metallicities (e.g. Belokurov and Kravtsov, 2022; Semenov et al., 2024; Chandra et al., 2023). Note that, as discussed earlier, the MW itself appears to be somewhat of an outlier: a dominant stellar disc emerges in the Galaxy at the earliest possible epochs, i.e. around $53<z<$ 5, but at metallicities much above the metal-poor regime, i.e. [Fe/H]$>-1.3$. At higher redshifts, the Galaxy is in its kinematically-hot, bursty and messy Aurora state. Sestito et al. (2021) address the origin of the low-metallicity stars on prograde, low-$|z|$ orbits, and conclude that most of them must be accreted. It remains unclear what induced the net spin in the low-metallicity halo, is it inherent or acquired? Dillamore et al. (2023a) demonstrate that the Galactic bar can spin up intrinsically non-rotating stellar haloes to velocities similar to those observed. In their models, a halo spun-up by the bar shows a rotation curve declining with Galactocentric radius — this prediction can be tested with upcoming data.
@@ 5.1 Sagittarius and the Milky Way disc
The Sgr dwarf galaxy (Ibata et al., 1994) and its associated stellar stream (e.g. Newberg et al., 2002; Majewski et al., 2003; Belokurov et al., 2006) are an archetype of dwarf galaxy disruption. The stellar stream emanating from a known dwarf galaxy is a clear example of accretion in action, and studies of the stream have been used to measure the Galactic potential as well as the orbit and mass of the Sgr dwarf progenitor (e.g. Law and Majewski, 2010; Dierickx and Loeb, 2017; Gibbons et al., 2017; Fardal et al., 2019; Vasiliev et al., 2021). The mere fact that the progenitor is still intact points to a fairly recent accretion event (since redshift $z\sim 1$), and this has been confirmed by most modelling efforts.
The omniscience of the Sgr stream has been obvious for several years, most notably displayed in the famous ‘field-of-streams’ image from SDSS (Belokurov et al., 2006). However, the Gaia mission has allowed a more detailed examination of this well-known stream (Antoja et al., 2020; Ibata et al., 2020; Ramos et al., 2020; Cunningham et al., 2024). In addition to properties of the stream itself, the Gaia era has provided evidence that Sgr has actually perturbed the Galactic disc during its orbit around the galaxy. Now, this is not a new idea — several past works have shown that a Sgr-like dwarf could induce perturbations to the disc (e.g. Dehnen, 1998; Ibata and Razoumov, 1998; Quillen et al., 2009; Laporte et al., 2018), and there has been evidence of disc perturbations from pre-Gaia data (e.g. Gómez et al., 2012; Widrow et al., 2012; Carlin et al., 2013; Schönrich and Dehnen, 2018). However, Gaia provided unambiguous evidence that the galaxy is out of equilibrium, and is currently undergoing a phase-mixing process from an out-of-equilibrium state. The evidence for this phase-mixing comes from the prominent spiral feature in position and velocity discovered by Antoja et al. (2018) (see Fig. 21). This so-called Gaia phase-space ‘Snail’ potentially arises from phase-mixing of past gravitational disturbances of the Milky Way disc. The spiral phase-space structures are studied in greater detail using Gaia DR3, where their chemical signatures can also be mapped thanks to the GSP-Spec catalogue derived from the RVS spectra (Gaia Collaboration et al., 2023).
@@ 5.2 A massive LMC
The most massive Milky Way satellites, the Large and Small Magellanic Clouds (LMC, SMC), have been known about for centuries. Visible to the naked eye in the Southern Hemisphere, their brightness is unmistakable. However, it is only in recent years that we have accepted the sheer massiveness of the Clouds, particularly the LMC. A simple abundance-matching argument already places the LMC at a significant mass: with a stellar mass of $2.7\times 10^{9}M_{\odot}$ (van der Marel et al., 2002), this roughly corresponds to a halo mass of $10^{11}M_{\odot}$ (Moster et al., 2013). A mass of $\sim 10^{11}M_{\odot}$ for the LMC also agrees with modelling efforts which aim to reproduce the significant 3D velocities of the Clouds and allow for LMC-SMC interactions (e.g. Besla et al., 2010; Kallivayalil et al., 2013). Most tidal models now advocate for a massive satellite (carrying its smaller sibling) that has rapidly joined the Milky Way in the past couple of Gyr (although Vasiliev 2024 show that the observations are also consistent with the LMC being on its second passage around the Milky Way). More recent works using Magellanic ‘satellite-of-satellite’ counts, or modelling of orbital streams in the presence of the LMC now agree that the LMC must have a mass $\gtrsim 1\times 10^{11}M_{\odot}$ (e.g. Erkal et al., 2019; Erkal and Belokurov, 2020)
At the pericentre of the LMC ($\sim 50$ kpc), a mass of $\sim 10^{11}M_{\odot}$ is a significant fraction of the Milky Way halo mass out to that radius (roughly $\sim 1/3$). Gómez et al. (2015) show that a massive LMC at this proximity can significantly displace the centre-of-mass (COM) of the Milky Way. Thus, this substantial perturber in our Galaxy cannot be ignored, and the scale of the damage inflicted by the LMC is now starting to be revealed by the Gaia mission.
@@ 5.2.1 LMC wake
Figure 22: Left: A model of the LMC-induced dynamical friction (transient) wake and collective response in the Milky Way halo shown in the Galactocentric YZ plane (Garavito-Camargo et al., 2021). The colorbar shows the density contrast (light = overdense, dark = underdense). The transient wake follows the past orbit of the LMC (indicated by the red line). The collective response appears predominantly to the North of the Galactic disc (central blue ellipse). Right: Distribution of K giant stars in Milky Way halo at 60 kpc $<R_{\rm gal}<100$ kpc in Galactic coordinates (Conroy et al., 2021). The overdensity in the North and South-West are attributed to the collective response and local wake, respectively. [Adapted from Garavito-Camargo et al. (2021); Conroy et al. (2021).]
@@ 5.2.2 The Milky Way-LMC system
The appreciation of the LMC’s impact on the Milky Way has spurned a new era of dynamical modelling: we now need to consider the ‘Milky Way-LMC system’ (or even the ‘Milky Way-LMC-Sgr system’) rather than just a Milky Way host halo. The fundamental assumptions that govern many modelling techniques, such as equilibrium, static potentials, spherical or axisymmetric symmetry, etc, are no longer valid. What are the implications of this for the many decades of work aiming to measure the total mass of the Galaxy using halo tracer populations (e.g. Wilkinson and Evans, 1999; Xue et al., 2008; Deason et al., 2012; Callingham et al., 2019; Eadie and Jurić, 2019)? A cautionary tale was spun by Erkal et al. (2021) who showed that mass estimates based on equilibrium modelling of tracers that ignore the presence of the LMC are typically biased high, and could even lead to an overestimate of up to 50 percent! On a more positive note, Deason et al. (2021) showed that by allowing for a small perturbation to the velocities of the halo stars a mass profile can be uncovered in a Milky Way-LMC potential. However, this depends on the radial range covered by the data and the region of the sky probed, as these authors focused on a region covered by SDSS where the predicted velocity gradients in the halo are small. Deason et al. (2021) also argued that systematic effects due to substructures, such as shells and clouds, can cause an even bigger effect than the LMC on mass-modelling efforts. Thus, one could argue that despite increasing sample sizes of multi-dimensional data, our aim to measure the Milky Way potential accurately, at least using classical methods, is perhaps waning.
@@
It is worth remarking in the post-Gaia era that we likely need to reassess what the ‘total mass of the Milky Way’ even means. Particularly in light of the messiness caused by the LMC and, to a lesser extent, Sgr. Detailed modelling of the Milky Way-LMC-Sgr system (Vasiliev et al., 2021) has indicated that considering these major components together can prove vital. Indeed, recent stream-modelling techniques have shown that the LMC is required in order to produce stream tracks that are consistent with the data (e.g. Erkal et al., 2019; Vasiliev et al., 2021). However, while tailored models are valuable there are still systematic effects revealed from cosmological simulations of stellar haloes that need to be taken into account in modelling efforts. How do we reconcile these two approaches? Perhaps the most promising avenue is the use of ‘Basis-Function-Expansion’ (BFE) methods (Lowing et al., 2011; Sanders et al., 2020; Garavito-Camargo et al., 2021; Petersen et al., 2022). Here, a time-evolving potential from an $N$-body simulation is approximated using a small number of basis functions. Thus, at a much lower cost than the original simulation, subhalo orbits can be explored in this potential without the need to continually re-simulate the host. In principle, these methods could be used to include the cosmological context of subhalo accretion, whilst retaining the ‘realistic’ potential of the Milky Way-LMC system. For example, Bayesian inference could be applied to a large suite of such models in order to infer global properties of the Galactic potential from the phase-space distribution of halo stars. The BFE models could even provide the framework to combine different dynamical modelling techniques, either from multiple streams and/or halo stars. With the exquisite data at hand, which will only increase in number and quality in the coming years, we clearly need to develop more sophisticated modelling approaches in order to keep pace with the data.
@@ 5.3 Census of accreted substructures
Figure 23: Dynamical substructure identified by Dodd et al. (2023) in the local stellar halo using the Gaia DR3 data. Top: Projections of the energy $E$ and angular momentum $L$ space. Bottom: Projections of the velocity space in cylindrical polars. [Reproduced from Dodd et al. (2023)].
@@
Figure 24: Streams discovered using STREAMFINDER algorithm, colour-coded according to their metallicity. [Reproduced from Ibata et al. (2024)].
@@
The identification by Helmi et al. (1999) of a handful of stars belonging to a stellar stream passing through the Solar neighourhood is a good example of the type of archaeological experiment that had appeared trailblazing and challenging before Gaia, but has been made effortless and almost mundane thanks to Gaia today. Since the pioneering study of Helmi et al. (1999) the methods to detect substructure have been sharpened (see e.g. Helmi and de Zeeuw, 2000; Harding et al., 2001; Knebe et al., 2005; Meza et al., 2005; Brown et al., 2005; Bullock and Johnston, 2005; Arifyanto and Fuchs, 2006; McMillan and Binney, 2008; Johnston et al., 2008; Gómez et al., 2010; Bovy et al., 2011; Sanderson et al., 2015; Lisanti et al., 2015; Antoja et al., 2015) whilst the available datasets have grown in size and richness (e.g. Yanny et al., 2000; Vivas et al., 2001; Newberg et al., 2002; Majewski et al., 2003; Clewley and Kinman, 2006; Kepley et al., 2007; Klement et al., 2008; Morrison et al., 2009; Klement et al., 2009; Schlaufman et al., 2009; Smith et al., 2009; Xue et al., 2011; Re Fiorentin et al., 2015; Janesh et al., 2016) but ultimately, the community had to wait for the arrival of the Gaia data to reveal the true wealth of accreted debris. This is particularly true for merger remnants that do not appear spatially coherent, although the detection of long, narrow streams has also evolved to a new level of sophistication thanks to the Gaia data (see e.g. Malhan and Ibata, 2018; Malhan et al., 2018; Ibata et al., 2019, 2024).
@@
Figure 25: Left: Current (black) and initial (red) Globular Cluster masses as a function of their orbital size. Assuming that the GC initial mass function is not a strong function of Galacto-centric radius, a large number of low-mass GC with apocentres inside $\sim$10 kpc appears to have been destroyed by the Galaxy. Middle and Right columns: Current (blue) and initial (red) mass functions of the Milky Way GCs in three different bins of Galacto-centric radius. Two different parametric forms for the initial mass functions are shown: power-law (middle) and log-normal (right). The integral of the difference between the current and the initial mass functions gives an estimate of the total number of GCs destroyed. [Adapted from Baumgardt et al. (2019)].
@@
Out of the box, the view on the halo substructure with Gaia DR1 appears underwhelming due to the minuscule number of halo stars in the shallow TGAS catalogue (Helmi et al., 2017; Myeong et al., 2017). Using the deeper SDSS-Gaia instead (see also Section 2.1), shows the promise of the modern space astrometry: the halo in the Solar neighourhood is riddled with pieces of accretion events (see Myeong et al., 2018a). This study uncovers $\lesssim 10$ new clumps in velocity space, of which at least one, namely S2, had been known before (Helmi et al., 1999). However, the integrals of motion space are better suited for substructure detection because phase-mixed debris can be smeared into arcs in velocity space; worse still, multiple wraps of the same stream might appear as distinct clumps. The immediate and clear improvement is demonstrated by Myeong et al. (2018b) who use SDSS-Gaia DR1 and translate into the action space, detecting a factor of $\sim 3$ more distinct sub-structures compared to their previous analysis of the same data (see also Wu et al., 2022, for a recent discussion of the pragmatic advantages of using action space). Once the Gaia DR2 was delivered, the hunt for the nearby halo merger remnants begun in earnest (Koppelman et al., 2018; Roederer et al., 2018; Koppelman et al., 2019; O’Hare et al., 2020; Borsato et al., 2020; Necib et al., 2020; Limberg et al., 2021; Lövdal et al., 2022; Tenachi et al., 2022; Ruiz-Lara et al., 2022; Malhan, 2022; Viswanathan et al., 2023; Dodd et al., 2023; Ye et al., 2024).